OpenStudy (anonymous):

help please 1 question

OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (−5, 5) and a directrix of y = -1. f(x) = −one twelfth (x − 5)2 + 2 f(x) = one twelfth (x − 5)2 + 2 f(x) = −one twelfth (x + 5)2 + 2 f(x) = one twelfth (x + 5)2 + 2

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

a actually went over this same question earlier today

OpenStudy (amistre64):

what are your thoughts on it so far?

OpenStudy (amistre64):

there are a few methods to approach this, so what you know is more important than what i know

OpenStudy (anonymous):

um well i dont really have an idea on where to start with this..can you you give me a little more to work wit sry

OpenStudy (amistre64):

parabolas have a defined structure, and the lessons/material in your course should cover these things. if you have no idea what to do at the moment, then you might want to review your material. if you had some idea about this i could help you iron out the rough spots.

OpenStudy (anonymous):

well i look at the lesson again but thx for try :)

OpenStudy (amistre64):

do you know a vertex form of a parabola? thats what the options are presented in where does a vertex fit with regards to a focus and directrix? how does the direction a parabola open effect the format?

OpenStudy (amistre64):

are you able to draw a rough sketch and label the information that is provided?

OpenStudy (amistre64):

stuff like this would be useful

OpenStudy (anonymous):

ok il look into the stuff you pointed out..thxman

OpenStudy (amistre64):

http://openstudy.com/users/amistre64#/updates/5436a767e4b09f2026ed562c and this is the same question as yours that i covered earlier that would give you pointers.

OpenStudy (anonymous):

kk thx

OpenStudy (amistre64):

yw